### v2

# A universal convergence theorem for primal-dual penalty and augmented Lagrangian methods

Abstract
We present a so-called universal convergence theorem for inexact primal-dual penalty and augmented Lagrangian methods that can be applied to a large number of such methods and reduces their convergence analysis to verification of some simple conditions on sequences generated by these methods. If these conditions are verified, then both primal and dual convergence follow directly from the universal convergence theorem. This theorem allows one not only to derive standard convergence theorems for many existing primal-dual penalty and augmented Lagrangian methods in a unified and straightforward manner, but also to strengthen and generalize some of these theorems. In particular, we show how with the use of the universal convergence theorem one can significantly improve some existing results on convergence of a primal-dual rounded weighted $\ell_{1}$-penalty method, an augmented Lagrangian method for cone constrained optimization, and some other primal-dual methods.
1 Introduction
Design and convergence analysis of primal-dual penalty and augmented Lagrangian methods have been active areas of research in constrained optimization for many years [1, 25, 24, 37, 4, 26]. Such methods, apart from being popular algorithms for solving various convex programming problems [31, 36, 21, 30, 39, 12], have also found applications in constrained global optimization [2, 11]. Furthermore, in the nonconvex case primal-dual methods combined with local search unconstrained optimization algorithms instead of the global ones sometimes significantly outperform specialized nonlinear programming solvers in computation time and/or quality of computed local minimizers (see the results of numerical experiments in [9, 17]).
Although convergence analysis of primal-dual penalty and augmented Lagrangian methods is a well-established topic in constrained optimization, little to no attention has been paid to uncovering general principles on which such analysis is based. If one carefully reads numerous papers dealing with convergence of primal-dual methods, one will quickly notice that very similar arguments are used in all these paper and very similar results are proved again and again for each new primal-dual method and each new type of constrained optimization problems (for example, compare convergence analysis from [3, 11, 9, 26, 38]). Some attempts to unify convergence analysis of augmented Lagrangian methods for inequality constrained problems were made in [25, 24, 37]. However, even in this specific case no real unification was achieved, since neither of the results of one of these papers can be derived from the results of the other two due to the use of different augmented Lagrangians and slightly different, non-interchangable assumptions, despite the fact that papers [25, 24, 37] contain very similar results that are based on very similar proofs.
The main goal of this paper is to present a so-called universal convergence theorem for primal-dual penalty and augmented Lagrangian methods that (i) uncovers general principles on which convergence analysis of such methods is based and (ii) can be applied to the vast majority of existing primal-dual methods for various types of constrained optimization problems, including nonlinear programming problems, cone constrained optimization problems, etc. To this end, we present a general duality scheme and an inexact primal-dual method based on this scheme that encompasses many existing primal-dual penalty and augmented Lagrangian methods as particular cases. Then we prove a general convergence theorem for this inexact primal-dual method that we call the universal convergence theorem. This theorem reduces the proof of primal and dual convergence to verification of some simple conditions on sequences generated by the method, the main one of which is the convergence of the infeasibility measure to zero. Verification of these conditions is at the core of the proofs of most existing convergence theorems for primal-dual methods. When these conditions are verified, the same standard arguments are used to prove primal and/or dual convergence. The universal convergence theorem eliminates the need to repeat these arguments.
In order to demonstrate the strengths of the universal convergence theorem and underline its universal nature, we consider its applications to convergence analysis of four different primal-dual penalty/augmented Lagrangian methods: (i) a simple primal-dual penalty method, (ii) a primal-dual rounded weighted $\ell_{1}$-penalty method proposed in [9], (iii) an extension of the classical augmented Lagrangian method based on the Hestenes-Powell-Rockafellar augmented Lagrangian [3] to the case of general cone constrained optimization problems and, finally, (iv) a multiplier algorithm based on the so-called P-type Augmented Lagrangian Method (PALM) from [37]. We show that with the use of the universal convergence theorem one can not only easily recover existing convergence theorems for these methods, but also significantly strengthen and generalize them.
The paper is organized as follows. A general duality scheme that allows one to describe many existing primal-dual penalty and augmented Lagrangian methods in a unified manner is presented in Section 2. Section 3 is devoted to the universal convergence theorem for inexact primal-dual methods, while Section 4 contains applications of this theorem to convergence analysis of several different penalty and augmented Lagrangian methods. Namely, in Subsection 4.1 we consider a simple primal-dual penalty method, in Subsection 4.2 a primal-dual rounded weighted $\ell_{1}$-penalty method, in Subsection 4.3 an augmented Lagrangian method for general cone constrained optimization problems, and in Subsection 4.4 a multiplier algorithm based on PALM.
2 General duality scheme
Let $(X,d_{X})$ be a metric space, $Y$ be a normed space, $Q\subset X$ be a nonempty closed set, $f\colon X\to\mathbb{R}\cup\{+\infty\}$ be a given function, and $\Phi\colon X\rightrightarrows Y$ be some multifunction. Throughout the first part of this article we study primal-dual methods for solving the following constrained optimization problem:
[…]
We assume that the feasible region of this problem is nonempty and its optimal value, denoted by $f_{*}$, is finite.
Let us introduce a general duality scheme for the problem $(\mathcal{P})$ that can be used to derive and analyse many existing primal-dual optimization methods in a unified manner. Let $(M,d_{M})$ be a metric space of parameters and suppose that some function $\varphi\colon X\times M\to(-\infty,+\infty]$ is given. We formally define a merit function for the problem $(\mathcal{P})$ as follows:
$F(x,\mu):=f(x)+\varphi(x,\mu)\quad x\in X,\>\mu\in M.$
Meanings of the function $\varphi$ and parameter $\mu$ depend on the context. If one wants to consider a penalty method, then $\mu\geq 0$ is a penalty parameter and one can define, for example, $\varphi(x,\mu)=\mu\operatorname{dist}(0,\Phi(x))$. If one consideres an augmented Lagrangian method, then one can set $\mu=(\lambda,c)$, where $\lambda$ is a multiplier and $c$ is a penalty parameter. In this case the function $\varphi(x,\mu)$ can be defined in many different ways. For example, if $\Phi$ is single-valued and $Y$ is a Hilbert space, then one can define $\varphi(x,\mu)=\langle\lambda,\Phi(x)\rangle+0.5c\|\Phi(x)\|^{2}$, where $\mu=(\lambda,c)$, $\lambda\in Y$, and $\langle\cdot,\cdot\rangle$ is the inner product.
In the general case, $\mu$ is simply a parameter of the merit function $F$, while $\varphi\colon X\times M\to(-\infty,+\infty]$ is a given function. No assumptions on this function are needed to formally define a duality scheme and a primal-dual method based on the merit function $F$. Some assumptions on the function $\varphi$ will appear later as a part of convergence analysis of this method.
The dual problem (associated with the merit function $F$) is defined in the following way:
[…]
where
$\Theta(\mu):=\inf_{x\in Q}F(x,\mu),\quad\mu\in M$
is the dual function (associated with the merit function $F$). The optimal value of the dual problem is denoted by $\Theta_{*}$. The following standard definition describes how the optimal values of the primal and dual problems should be related to each other in order to make primal-dual methods based on the proposed duality scheme tractable.
Definition 1.
One says that the weak duality holds (for the pair of problems $(\mathcal{P})$ and $(\mathcal{D})$), if $\Theta_{*}\leq f_{*}$. One says that the strong duality holds (or that the zero duality gap property holds true for the merit function $F$), if $\Theta_{*}=f_{*}$.
Let us point out a natural assumption that ensures the weak duality.
Lemma 1.
Suppose that for any $\mu\in M$ and any feasible point $x$ of the problem $(\mathcal{P})$ one has $\varphi(x,\mu)\leq 0$. Then the weak duality holds for the pair of problems $(\mathcal{P})$ and $(\mathcal{D})$.
Proof.
Denote by $\Omega$ the feasible region of the problem $(\mathcal{P})$. By our assumption for any $x\in\Omega$ and $\mu\in M$ one has $F(x,\mu)\leq f(x)$. Therefore
$\Theta(\mu):=\inf_{x\in Q}F(x,\mu)\leq\inf_{x\in\Omega}F(x,\mu)\leq\inf_{x\in\Omega}f(x)=f_{*}\quad\forall\mu\in M,$
which implies that $\Theta_{*}\leq f_{*}$. ∎
3 Inexact primal-dual method and its convergence analysis
Let us now turn to analysis of primal-dual methods for solving the problem $(\mathcal{P})$ that are based on the merit function $F$. A large class of existing primal-dual penalty and augmented Lagrangian methods can be described with the use of the following inexact primal-dual method given in Algorithm 1.
Initialization. Choose an initial value of the parameter $\mu_{0}\in M$ and a bounded above sequence of tolerances $\{\varepsilon_{n}\}\subset[0,+\infty)$. Put $n=0$.
Step 1. Solution of subproblem. Find an $\varepsilon_{n}$-optimal solution $x_{n}$ of the problem
$\min_{x}\>F(x,\mu_{n})\quad\text{subject to}\quad x\in Q,$
that is, find $x_{n}\in Q$ such that $F(x_{n},\mu_{n})\leq F(x,\mu_{n})+\varepsilon_{n}$ for all $x\in Q$.
Step 2. Update of parameters. Choose some $\mu_{n+1}\in M$, increment $n$, and go to Step 1.
Algorithm 1 Inexact primal-dual method
Throughout this article we impose the following assumption on sequences generated by primal-dual methods that is needed to ensure that the primal sequence $\{x_{n}\}$ is correctly defined.
Assumption 1.
For any $n\in\mathbb{N}$ the function $F(\cdot,\mu_{n})$ is bounded below on $Q$.
It should be noted that we do not assume that $\varepsilon_{n}\to 0$ as $n\to\infty$, that is, we do not assume that an approximate solution of the subproblem
$\min_{x}\>F(x,\mu_{n})\quad\text{subject to}\quad x\in Q$ (1)
can be computed with arbitrarily high precision, as it is often done in the literature (see, e.g. [25, 37, 6, 4, 7]). The value $\varepsilon_{n}>0$ represents not how accurately one needs to solve problem (1), but rather how accurately one can solve it in a reasonable amount of time. The assumption that $\varepsilon_{n}\to 0$ as $n\to\infty$ is purely theoretical in nature and is needed only for better general understanding of asymptotic behaviour of sequences generated by primal-dual methods. That is why in the case when $\varepsilon_{n}\to 0$ as $n\to\infty$ we call Algorithm 1 the asymptotically exact primal-dual method.
Remark 1.
Throughout this article we are interested only in the convergence analysis of infinite sequences generated by primal-dual penalty and augmented Lagrangian methods. Therefore, we do not include stopping criteria in the descriptions of these methods and always assume that a method under consideration generates an infinite sequence of points whose convergence to (approximately) globally optimal solutions we are aiming to prove. Natural stopping criteria for these methods can be easily designed using the results of our convergence analysis.
Our aim is to prove a general convergence theorem for the inexact primal-dual method, which we call a universal convergence theorem. This theorem reduces convergence analysis of primal-dual methods to verification of some natural conditions. In the following section we will demonstrate that in many particular cases these conditions can be verified in a straightforward manner.
The fundamental assumption that one must impose on primal-dual methods is the decay of the infeasibility measure. If the infeasibility measure does not converge to zero, then the corresponding method fails to solve an optimization problem to which it was applied. In our case, the decay of the infeasibility measure can be expressed as an assumption that $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$, where $\operatorname{dist}(0,\Phi(x_{n}))=\inf_{y\in\Phi(x_{n})}\|y\|$ is the distance between 0 and $\Phi(x_{n})$. This condition along with an additional assumption on the sequence $\{\varphi(x_{n},\mu_{n})\}$ are, roughly speaking, enough to draw multiple conclusions about the behaviour of the sequence $\{(x_{n},\mu_{n})\}$ generated by the inexact primal-dual method. To conveniently formulate them, recall that the function
$\beta(y)=\inf\Big{\{}f(x)\Bigm{|}x\in Q\colon 0\in\Phi(x)-y\Big{\}}$
is called the optimal value (perturbation) function for the problem $(\mathcal{P})$. In addition, denote $\varepsilon_{*}=\limsup_{n\to\infty}\varepsilon_{n}$. Note that $0\leq\varepsilon_{*}<+\infty$, since by our assumption the sequence $\{\varepsilon_{n}\}$ is bounded above.
Lemma 2.
Let the weak duality hold and $\{(x_{n},\mu_{n})\}$ be the sequence generated by the inexact primal-dual method. Suppose that
$\lim_{n\to\infty}\operatorname{dist}(0,\Phi(x_{n}))=0,\quad\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0.$ (2)
Then the following inequalities hold true:
$\min\Big{\{}f_{*},\liminf_{y\to 0}\beta(y)\Big{\}}\leq\liminf_{n\to\infty}f(x_{n})$ $\leq\limsup_{n\to\infty}f(x_{n})\leq f_{*}+\varepsilon_{*},$ (3)
$\min\Big{\{}f_{*},\liminf_{y\to 0}\beta(y)\Big{\}}\leq\liminf_{n\to\infty}F(x_{n},\mu_{n})$ $\leq\limsup_{n\to\infty}F(x_{n},\mu_{n})\leq f_{*}+\varepsilon_{*},$ (4)
$\min\Big{\{}f_{*},\liminf_{y\to 0}\beta(y)\Big{\}}-\varepsilon_{*}\leq\liminf_{n\to\infty}\Theta(\mu_{n})$ $\leq\limsup_{n\to\infty}\Theta(\mu_{n})\leq f_{*}.$ (5)
Proof.
Part 1. Upper estimates. Since the weak duality holds, by the definition of $x_{n}$ one has
$f_{*}\geq\Theta_{*}\geq\Theta(\mu_{n})\geq F(x_{n},\mu_{n})-\varepsilon_{n}=f(x_{n})+\varphi(x_{n},\mu_{n})-\varepsilon_{n}\quad\forall n\in\mathbb{N},$
which yields
$\limsup_{n\to\infty}\Theta(\mu_{n})\leq f_{*},\quad\limsup_{n\to\infty}F(x_{n},\mu_{n})\leq f_{*}+\varepsilon_{*}$
and
$\limsup_{n\to\infty}f(x_{n})$ $\leq\limsup_{n\to\infty}\big{(}f_{*}+\varepsilon_{n}-\varphi(x_{n},\mu_{n})\big{)}$
$\leq f_{*}+\limsup_{n\to\infty}\varepsilon_{n}+\limsup_{n\to\infty}(-\varphi(x_{n},\mu_{n}))$
$=f_{*}+\varepsilon_{*}-\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\leq f_{*}+\varepsilon_{*}.$
that is, the upper estimates in (3)–(5) are satisfied.
Part 2. Lower estimates. Choose any subsequence $\{x_{n_{k}}\}$ such that
$\lim_{k\to\infty}f(x_{n_{k}})=\liminf_{n\to\infty}f(x_{n}).$
At least one such subsequence exists by the definition of lower limit. Let us consider two cases.
Case I. Suppose that there exists a subsequence of the sequence $\{x_{n_{k}}\}$, which we denote again by $\{x_{n_{k}}\}$, that is feasible for the problem $(\mathcal{P})$. Then $f(x_{n_{k}})\geq f_{*}$ for all $k\in\mathbb{N}$ and
[…]
Case II. Suppose that a feasible subsequence of the sequence $\{x_{n_{k}}\}$ does not exist. Then without loss of generality one can suppose that every element of the sequence $\{x_{n_{k}}\}$ is infeasible. By the definition of optimal value function for any $k\in\mathbb{N}$ and $y_{k}\in Y$ such that $0\in\Phi(x_{n_{k}})-y_{k}$ one has $f(x_{n_{k}})\geq\beta(y_{k})$. Recall that $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$ by our assumption. Consequently, one can find a sequence $\{y_{k}\}\subset Y$ such that $0\in\Phi(x_{n_{k}})-y_{k}$ for all $k\in\mathbb{N}$ and $\|y_{k}\|\to 0$ as $k\to\infty$. Therefore
$\liminf_{n\to\infty}f(x_{n})=\lim_{k\to\infty}f(x_{n_{k}})\geq\liminf_{k\to\infty}\beta(y_{k})\geq\liminf_{y\to 0}\beta(y).$
Combining the two cases one can conclude that the lower estimate in (3) holds true. Bearing in mind the second condition in (2) one gets
$\liminf_{n\to\infty}F(x_{n},\mu_{n})=\liminf_{n\to\infty}\big{(}f(x_{n})+\varphi(x_{n},\mu_{n})\big{)}\geq\liminf_{n\to\infty}f(x_{n}),$
that is, the lower estimate in (4) is valid. Finally, by the definition of $x_{n}$ one has
$\liminf_{n\to\infty}\Theta(\mu_{n})\geq\liminf_{n\to\infty}(F(x_{n},\mu_{n})-\varepsilon_{n})\geq\liminf_{n\to\infty}F(x_{n},\mu_{n})-\varepsilon_{*},$
which along with the first inequality in (4) implies that the lower estimate in (5) holds true. ∎
One can strengthen the claim of the previous lemma with the use of an additional technical assumption on the sequence $\{\mu_{n}\}$/the function $\varphi$ that, as we will show below, can be easily verified in many particular cases.
Definition 2.
A sequence $\{\mu_{n}\}\subset M$ is said to satisfy the $\varphi$-convergence to zero condition, if there exists a sequence $\{t_{n}\}\subset(0,+\infty)$ such that for any sequence $\{z_{n}\}\subset Q$ satisfying the inequality $\operatorname{dist}(0,\Phi(z_{n}))\leq t_{n}$ for all $n\in\mathbb{N}$ one has $\limsup_{n\to\infty}\varphi(z_{n},\mu_{n})\leq 0$.
Corollary 1.
Let the assumptions of Lemma 2 hold true and suppose that the sequence $\{\mu_{n}\}$ satisfies the $\varphi$-convergence to zero condition. Then one has $\limsup_{n\to\infty}\varphi(x_{n},\mu_{n})\leq\varepsilon_{*}$ and the following inequalities hold true:
$\vartheta_{*}\leq\liminf_{n\to\infty}f(x_{n})$ $\leq\limsup_{n\to\infty}f(x_{n})\leq\vartheta_{*}+\varepsilon_{*},$ (6)
$\vartheta_{*}\leq\liminf_{n\to\infty}F(x_{n},\mu_{n})$ $\leq\limsup_{n\to\infty}F(x_{n},\mu_{n})\leq\vartheta_{*}+\varepsilon_{*},$ (7)
$\vartheta_{*}-\varepsilon_{*}\leq\liminf_{n\to\infty}\Theta(\mu_{n})$ $\leq\limsup_{n\to\infty}\Theta(\mu_{n})\leq\vartheta_{*}.$ (8)
where $\vartheta_{*}=\min\{f_{*},\liminf_{y\to 0}\beta(y)\}$.
Proof.
Denote $\beta_{*}=\liminf_{y\to 0}\beta(y)$. If $f_{*}\leq\beta_{*}$, then the claim of the corollary follows directly from Lemma 2. Therefore, suppose that $\beta_{*}<f_{*}$. Note that the lower estimates in (6)–(8) also follow from Lemma 2, which means that we need to prove only the upper estimates.
By the definition of lower limit there exists a sequence $\{y_{k}\}\subset Y$ such that $y_{k}\to 0$ and $\beta(y_{k})\to\beta_{*}$ as $k\to\infty$. Since the sequence $\{y_{k}\}$ converges to zero, there exists a subsequence $\{y_{k_{n}}\}$ such that $\|y_{k_{n}}\|\leq t_{n}$ for all $n\in\mathbb{N}$, where $t_{n}$ is from Definition 2. In turn, with the use of the definition of the optimal value function $\beta$ one can find a sequence $\{z_{n}\}\subset Q$ such that
$0\in\Phi(z_{n})-y_{k_{n}}\quad\forall n\in\mathbb{N},\quad\lim_{n\to\infty}f(z_{n})=\lim_{n\to\infty}\beta(y_{k_{n}}).$
Note that $\operatorname{dist}(0,\Phi(z_{n}))\leq\|y_{k_{n}}\|\leq t_{n}$ for all $n\in\mathbb{N}$. Hence by the $\varphi$-convergence to zero condition one has $\limsup_{n\to\infty}\varphi(z_{n},\mu_{n})\leq 0$ and
$\limsup_{n\to\infty}\Theta(\mu_{n})\leq\limsup_{n\to\infty}F(z_{n},\mu_{n})$ $\leq\limsup_{n\to\infty}f(z_{n})+\limsup_{n\to\infty}\varphi(z_{n},\mu_{n})$
$\leq\limsup_{n\to\infty}f(z_{n})=\lim_{n\to\infty}\beta(y_{k_{n}})=\beta_{*},$
which means that the upper estimate in (8) holds true.
Recall that $F(x_{n},\mu_{n})\leq\Theta(\mu_{n})+\varepsilon_{n}$ for all $n\in\mathbb{N}$ by the definition of $x_{n}$. Consequently, the upper estimate in (7) follows directly from the upper estimate in (8). With the use of this upper estimate and the second inequality in (2) one obtains
$\limsup_{n\to\infty}f(x_{n})$ $=\limsup_{n\to\infty}\Big{(}F(x_{n},\mu_{n})-\varphi(x_{n},\mu_{n})\Big{)}$
$\leq\limsup_{n\to\infty}F(x_{n},\mu_{n})+\limsup_{n\to\infty}\big{(}-\varphi(x_{n},\mu_{n})\big{)}$
$\leq\vartheta_{*}+\varepsilon_{*}-\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\leq\vartheta_{*}+\varepsilon_{*},$
that is, the upper estimate in (6) is satisfied. Finally, note that
$\limsup_{n\to\infty}\varphi(x_{n},\mu_{n})$ $=\limsup_{n\to\infty}\Big{(}F(x_{n},\mu_{n})-f(x_{n})\Big{)}$
$\leq\limsup_{n\to\infty}F(x_{n},\mu_{n})-\liminf_{n\to\infty}f(x_{n})\leq\vartheta_{*}+\varepsilon_{*}-\vartheta_{*}=\varepsilon_{*},$
which completes the proof of the corollary. ∎
With the use of Lemma 2 and Corollary 1 we can prove a general convergence theorem for the inexact primal-dual method (Algorithm 1), which we call the universal convergence theorem.
Theorem 1 (universal convergence theorem for inexact primal-dual methods).
Let the weak duality hold and $\{(x_{n},\mu_{n})\}$ be the sequence generated by the inexact primal-dual method. Suppose that
$\lim_{n\to\infty}\operatorname{dist}(0,\Phi(x_{n}))=0,\quad\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0.$
Then
$\underline{\vartheta}_{*}\leq\liminf_{n\to\infty}f(x_{n})$ $\leq\limsup_{n\to\infty}f(x_{n})\leq\overline{\vartheta}_{*}+\varepsilon_{*},$ (9)
$\underline{\vartheta}_{*}\leq\liminf_{n\to\infty}F(x_{n},\mu_{n})$ $\leq\limsup_{n\to\infty}F(x_{n},\mu_{n})\leq\overline{\vartheta}_{*}+\varepsilon_{*},$ (10)
$\underline{\vartheta}_{*}-\varepsilon_{*}\leq\liminf_{n\to\infty}\Theta(\mu_{n})$ $\leq\limsup_{n\to\infty}\Theta(\mu_{n})\leq\overline{\vartheta}_{*},$ (11)
where $\underline{\vartheta}_{*}=\min\{f_{*},\liminf_{y\to 0}\beta(y)\}$, $\overline{\vartheta}_{*}=\underline{\vartheta}_{*}$, if the sequence $\{\mu_{n}\}$ satisfies the $\varphi$-convergence to zero condition, and $\overline{\vartheta}_{*}=f_{*}$ otherwise.
In addition, if the optimal value function $\beta$ is lower semicontinuous at the origin, then $\underline{\vartheta}_{*}=\overline{\vartheta}_{*}=f_{*}$, $\limsup_{n\to\infty}\varphi(x_{n},\mu_{n})\leq\varepsilon_{*}$, and the two following statements hold true:
  1. 1.
    if the functions $f$ and $\operatorname{dist}(0,\Phi(\cdot))$ are lsc on $Q$ and the set-valued map $\Phi$ takes closed values on the set $Q$, then any limit point of the sequence $\{x_{n}\}$ (if such point exists) is an $\varepsilon_{*}$-optimal solution of the problem $(\mathcal{P})$;
  2. 2.
    if the function $\Theta$ is upper semicontinuous (usc), then any limit point of the sequence $\{\mu_{n}\}$ (if such point exists) is an $\varepsilon_{*}$-optimal solution of the dual problem $(\mathcal{D})$.
Proof.
Inequalities (9)–(11) follow directly from Lemma 2 and Corollary 1.
Suppose that $\beta$ is lsc at the origin. By the definition of optimal value function $\beta(0)=f_{*}$. Therefore, the lower semicontinuity assumption is equivalent to the condition $\min\{f_{*},\lim_{y\to 0}\beta(y)\}=f_{*}$, which yields $\underline{\vartheta}_{*}=\overline{\vartheta}_{*}=f_{*}$.
With the use of inequalities (10) and (9) one gets
$f_{*}+\varepsilon_{*}\geq\limsup_{n\to\infty}F(x_{n},\mu_{n})$ $\geq\liminf_{n\to\infty}f(x_{n})+\limsup_{n\to\infty}\varphi(x_{n},\mu_{n})$
$\geq f_{*}+\limsup_{n\to\infty}\varphi(x_{n},\mu_{n}),$
which implies that $\limsup_{n\to\infty}\varphi(x_{n},\mu_{n})\leq\varepsilon_{*}$.
Let $x_{*}$ be a limit point of the sequence $\{x_{n}\}$, that is, some subsequence $\{x_{n_{k}}\}$ converges to $x_{*}$. Note that $x_{*}\in Q$, since $Q$ is a closed set. In addition, $\operatorname{dist}(0,\Phi(x_{*}))=0$ due to the facts that $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$ and the function $\operatorname{dist}(0,\Phi(\cdot))$ is lsc. Hence bearing in mind the fact that the set $\Phi(x_{*})$ is closed by our assumption one can conclude that $x_{*}$ is a feasible point of the problem $(\mathcal{P})$. Moreover, by the upper estimate in (9) and the lower semicontinuity of $f$ one gets that $f(x_{*})\leq f_{*}+\varepsilon_{*}$, that is, $x_{*}$ is an $\varepsilon_{*}$-optimal solution of the problem $(\mathcal{P})$. The statement about limit points of the sequence $\{\mu_{n}\}$ is proved in the same way. ∎
For the sake of completeness, let us also point out how the previous theorem can be strengthened in the case of the asymptotically exact primal-dual method, that is, in the case when $\varepsilon_{n}\to 0$ as $n\to\infty$ (or, equivalently, $\varepsilon_{*}=0$).
Theorem 2 (universal convergence theorem for asymptotically exact primal-dual methods).
Let the weak duality hold, the optimal value function $\beta$ be lsc at the origin, and $\{(x_{n},\mu_{n})\}$ be the sequence generated by the asymptotically exact primal-dual method. Suppose that
$\lim_{n\to\infty}\operatorname{dist}(0,\Phi(x_{n}))=0,\quad\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0.$
Then the strong duality holds, $\lim_{n\to\infty}\varphi(x_{n},\mu_{n})=0$, and
$f_{*}=\lim_{n\to\infty}f(x_{n})=\lim_{n\to\infty}F(x_{n},\mu_{n})=\lim_{n\to\infty}\Theta(\mu_{n})=\Theta_{*}.$ (12)
In addition, the following statements hold true:
  1. 1.
    if the functions $f$ and $\operatorname{dist}(0,\Phi(\cdot))$ are lsc on $Q$ and the set-valued map $\Phi$ takes closed values on the set $Q$, then any limit point of the sequence $\{x_{n}\}$ (if such point exists) is a globally optimal solution of the problem $(\mathcal{P})$;
  2. 2.
    if the function $\Theta$ is usc, then any limit point of the sequence $\{\mu_{n}\}$ (if such point exists) is a globally optimal solution of the dual problem $(\mathcal{D})$.
Proof.
Note that by the weak duality and the definition of $\Theta_{*}$ one has
$\limsup_{n\to\infty}\Theta(\mu_{n})\leq\Theta_{*}\leq f_{*}.$
Bearing in mind these inequalities and putting $\varepsilon_{*}=0$ in Theorem 1 one obtains the required result. ∎
Remark 2.
In many particular cases the assumption that the optimal value function $\beta$ is lsc at the origin is, in fact, equivalent to the strong duality [33, 18]. Therefore, in these cases the previous theorem can be restated as follows. If the strong duality and conditions (2) hold, then equalities (12) are satisfied.
In the case when either $X$ or $M$ is a reflexive normed space (in particular, finite dimensional normed space) one can prove slightly stronger results. For the sake of shortness we formulate these results only for the asymptotically exact primal-dual method.
Corollary 2.
Let $X$ be a reflexive normed space, the functions $f$ and $\operatorname{dist}(0,\Phi(\cdot))$ be weakly sequentially lsc on $Q$, the set-valued map $\Phi$ take closed values on the set $Q$, the set $Q$ be weakly sequentially closed (in particular, convex), and the assumptions of Theorem 2 be valid. Then the following statements hold true:
  1. 1.
    any weakly limit point of the sequence $\{x_{n}\}$ (if such point exists) is a globally optimal solution of the problem $(\mathcal{P})$;
  2. 2.
    if the sequence $\{x_{n}\}$ is bounded and a globally optimal solution of the problem $(\mathcal{P})$ is unique, then the sequence $\{x_{n}\}$ weakly converges to this globally optimal solution.
Proof.
The validity of the first statement is proved in exactly the same way as the validity of the analogous statement from Theorem 1.
Let us prove the second statement. Indeed, suppose by contradiction that the sequence $\{x_{n}\}$ does not weakly converge to unique globally optimal solution $x_{*}$ of the problem $(\mathcal{P})$. Then there exists a subsequence $\{x_{n_{k}}\}$ and a neighbourhood $U(x_{*})$ of $x_{*}$ in the weak topology such that $x_{n_{k}}\notin U(x_{*})$ for all $k\in\mathbb{N}$.
By our assumption the sequence $\{x_{n_{k}}\}$ is bounded. Hence taking into account the fact that the space $X$ is reflexive one can conclude that there exists a subsequence $\{x_{n_{k_{\ell}}}\}$ that weakly converges to some point $z_{*}$. By the first statement of the corollary the point $z_{*}$ is a globally optimal solution of the problem $(\mathcal{P})$. Thus, $z_{*}=x_{*}$, which is impossible, since $\{x_{n_{k_{\ell}}}\}$ weakly converges to $z_{*}$, but $x_{n_{k}}\notin U(x_{*})$ for all $k\in\mathbb{N}$. ∎
Corollary 3.
Let $M$ be a reflexive normed space, the function $\Theta$ weakly sequentially usc, and the assumptions of Theorem 2 be valid. Then the following statements hold true:
  1. 1.
    any weakly limit point of the sequence $\{\mu_{n}\}$ (if such point exists) is a globally optimal solution of the dual problem $(\mathcal{D})$;
  2. 2.
    if the sequence $\{\mu_{n}\}$ is bounded and a globally optimal solution of the dual problem is unique, then the sequence $\{\mu_{n}\}$ weakly converges to this globally optimal solution.
Remark 3.
All results of this section can be slightly generalized in the following way. Namely, one can easily verify that they remain to hold true, if instead of the weak duality one simply assumes that $\limsup_{n\to\infty}\Theta(\mu_{n})\leq f_{*}$.
4 Applications
The goal of this section is to show that many existing convergence theorems for various penalty and augmented Lagrangian methods can be easily proved with the use of the universal convergence theorem for the inexact primal-dual method (Theorem 1) in a unified manner.
4.1 Primal-dual penalty method
Let us first consider a simple illustrative example of a primal-dual penalty method for the problem $(\mathcal{P})$. Let $\mathbb{R}_{+}=[0,+\infty)$ and $\omega\colon\mathbb{R}_{+}\to\mathbb{R}_{+}\cup\{$$\infty\}$ be a nondecreasing function such that $\operatorname{dom}\omega=[0,\tau)$ for some $\tau\in(0,+\infty]$, $\omega(0)=0$, the function $\omega$ is strictly increasing and continuous on $0,\tau)$, and $\omega(t)\to+\infty$ as $t\to\tau$. Introduce the penalty function for the problem $(\mathcal{P})$ as follows:
$F(x,c)=f(x)+c\omega\big{(}\operatorname{dist}(0,\Phi(x))\big{)}\quad x\in X,\>c\geq 0.$ (13)
Here $c\geq 0$ is the penalty parameter and $\omega(+\infty)=+\infty$ by definition. In this case $M=[0,+\infty)$, $\mu=c$, and $\varphi(x,\mu)=\mu\omega(\operatorname{dist}(0,\Phi(x)))$. The dual function
$\Theta(c)=\inf_{x\in Q}\Big{(}f(x)+c\omega\big{(}\operatorname{dist}(0,\Phi(x))\big{)}\Big{)}\quad\forall c\geq 0.$
in this case is obviously concave and usc.
If a point $x_{n}\in Q$ is an $\varepsilon_{n}$-optimal solution of the problem
$\min_{x}\enspace f(x)+c_{n}\omega\big{(}\operatorname{dist}(0,\Phi(x))\big{)}\quad\text{subject to}\quad x\in Q$ (14)
for some $c_{n}>0$ and $\varepsilon_{n}>0$, then by the definitions of $\Theta$ and $x_{n}$ for any $c\geq 0$ one has
$\Theta(c)-\Theta(c_{n})$ $\leq f(x_{n})+c\omega\big{(}\operatorname{dist}(0,\Phi(x_{n}))\big{)}-f(x_{n})-c_{n}\omega\big{(}\operatorname{dist}(0,\Phi(x_{n}))\big{)}+\varepsilon_{n}$
$=\omega\big{(}\operatorname{dist}(0,\Phi(x_{n}))\big{)}\big{(}c-c_{n}\big{)}+\varepsilon_{n}.$
Thus, $p=\omega(\operatorname{dist}(0,\Phi(x_{n})))$ is an $\varepsilon_{n}$-supergradient of the concave function $\Theta$, and one can use $p$ in order to define a subgradient-type step as an update of the penalty parameter:
$c_{n+1}=c_{n}+s_{n}\big{[}\omega(\operatorname{dist}(0,\Phi(x_{n})))+\delta_{n}\big{]},\quad s_{n}>0,\quad\delta_{n}\geq 0.$
Here $s_{n}$ is the step length, while $\delta_{n}\geq 0$ is a correction. Based on this penalty parameter update, we can define the following primal-dual penalty method given in Algorithm 2. Let us note that $p=\omega(\operatorname{dist}(0,\Phi(x_{n})))$ is necessarily finite, since the optimal value of problem (14) does not exceed $f_{*}$ for any $c_{n}\geq 0$.
Initialization. Choose an initial value of the penalty parameter $c_{0}\geq 0$ and a bounded above sequence of tolerances $\{\varepsilon_{n}\}\subset[0,+\infty)$. Put $n=0$.
Step 1. Solution of subproblem. Find an $\varepsilon_{n}$-optimal solution $x_{n}$ of the problem
$\min_{x}\>f(x)+c_{n}\omega\big{(}\operatorname{dist}(0,\Phi(x))\big{)}\quad\text{subject to}\quad x\in Q.$
Step 2. Penalty parameter update. Choose some $s_{n}>0$, $\delta_{n}\geq 0$, and define
$c_{n+1}=c_{n}+s_{n}\big{[}\omega\big{(}\operatorname{dist}(0,\Phi(x_{n}))\big{)}+\delta_{n}\big{]}.$
Increment $n$ and go to Step 1.
Algorithm 2 Primal-dual penalty method
Let us show how one can easily prove convergence of the primal-dual penalty method with the use of the universal convergence theorem. For the sake of shortness we will only consider the general non-reflexive case and will not show how Corollary 2 can be reformulated for the primal-dual penalty method.
As above, $\varepsilon_{*}=\limsup_{n\to\infty}\varepsilon_{n}$. Recall also that the penalty function $F$ is called globally exact [13, 41, 14], if there exists $c_{n}>0$ such that problem (14) has the same globally optimal solutions as the problem $(\mathcal{P})$. The infimum of all such $c_{n}$ is denoted by $c_{*}(F)$ and is called the least exact penalty parameter for the penalty function $F$.
Theorem 3.
Let the penalty function $F(\cdot,c_{0})$ be bounded below and the sequence $\{(x_{n},c_{n})\}$ be generated by the primal-dual penalty method. Suppose also that there exists $s_{*}>0$ such that $s_{n}\geq s_{*}$ for all $n\in\mathbb{N}$. Then $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$ and
$\min\{f_{*},\liminf_{y\to 0}\beta(y)\}$ $\leq\liminf_{n\to\infty}f(x_{n})$
$\leq\limsup_{n\to\infty}f(x_{n})\leq\min\{f_{*},\liminf_{y\to 0}\beta(y)\}+\varepsilon_{*}.$
In addition, if the optimal value function $\beta$ is lsc at the origin, then
$\limsup_{n\to\infty}c_{n}\omega\big{(}\operatorname{dist}(0,\Phi(x_{n}))\big{)}\leq\varepsilon_{*}$
and the following statements hold true:
  1. 1.
    if the functions $f$ and $\operatorname{dist}(0,\Phi(\cdot))$ are lsc on $Q$ and the set-valued map $\Phi$ takes closed values on the set $Q$, then any limit point of the sequence $\{x_{n}\}$ (if such point exists) is an $\varepsilon_{*}$-optimal solution of the problem $(\mathcal{P})$;
  2. 2.
    if $\varepsilon_{*}=0$ and the sequence $\{c_{n}\}$ is bounded, then the penalty function $F$ is globally exact and $c_{*}:=\lim_{n\to\infty}c_{n}\geq c_{*}(F)$.
Proof.
Since the sequence $\{c_{n}\}$ is obviously nondecreasing, the boundedness below of the function $F(\cdot,c_{0})$ guarantees that Assumption 1 is satisfied.
As was noted above, in the case under consideration one can define $\varphi(x,\mu)=\mu\omega(\operatorname{dist}(0,\Phi(x))$ for any $\mu\in M=\mathbb{R}_{+}$ and $x\in X$. This function is nonnegative and $\varphi(x,\mu)=0$ for any $\mu\in M$ and any feasible point $x$. Hence by Lemma 1 the weak duality holds and $\liminf_{n\to\infty}\varphi(x_{n},c_{n})\geq 0$. Furthermore, the sequence $\{c_{n}\}$ obviously satisfies the $\varphi$-convergence to zero condition with any sequence $\{t_{n}\}\subset(0,+\infty)$ such that $c_{n}\omega(t_{n})\to 0$ as $n\to\infty$ (at least one such sequence $\{t_{n}\}$ exists, since $\omega(0)=0$ and $\omega$ is continuous on some interval $0,\tau)$). Therefore, in order to apply the universal convergence theorem for the inexact primal-dual method it is sufficient to check that $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$.
Suppose by contradiction that $\operatorname{dist}(0,\Phi(x_{n}))$ does not converge to zero. Then the sequence $\omega(\operatorname{dist}(0,\Phi(x_{n}))$ does not converge to zero either and, therefore,
$\sum_{n=0}^{\infty}\omega\big{(}\operatorname{dist}(0,\Phi(x_{n}))\big{)}=+\infty.$ (15)
According to Step 2 of Algorithm 2 and our assumption on the sequence $\{s_{n}\}$ one has
$c_{n}=c_{0}+\sum_{i=0}^{n-1}\Big{(}s_{i}\omega\big{(}\operatorname{dist}(0,\Phi(x_{i}))\big{)}+\delta_{i}\Big{)}\geq c_{0}+s_{*}\sum_{i=0}^{n-1}\omega\big{(}\operatorname{dist}(0,\Phi(x_{i}))\big{)}$
for any $n\in\mathbb{N}$. Hence taking into account (15) one can conclude that $c_{n}\to\infty$ as $n\to\infty$.
On the other hand, note that
$F(x_{n},c_{n})$ $=F(x_{n},c_{0})+(c_{n}-c_{0})\omega\big{(}\operatorname{dist}(0,\Phi(x_{n})\big{)}$
$\geq\inf_{x\in Q}F(x,c_{0})+(c_{n}-c_{0})\omega\big{(}\operatorname{dist}(0,\Phi(x_{n})\big{)}$
for any $n\in\mathbb{N}$. Since the sequence $\{c_{n}\}$ increases unboundedly, while the sequence $\{\omega(\operatorname{dist}(0,\Phi(x_{n}))\}$ does not converge to zero, the inequality above implies that
$\limsup_{n\to\infty}F(x_{n},c_{n})=+\infty.$ (16)
However, by the weak duality one has
$F(x_{n},c_{n})\leq\inf_{x\in Q}F(x,c_{n})+\varepsilon_{n}\leq f_{*}+\varepsilon_{n},$
which contradicts (16). Thus, $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$ and all claims of the theorem, except for the claim about the sequence $\{c_{n}\}$, hold true by Theorem 1.
If $\varepsilon_{*}=0$ and the sequence $\{c_{n}\}$ is bounded, then its limit $c_{*}$ (which exists since the sequence $\{c_{n}\}$ is nondecreasing) is a globally optimal solution of the dual problem and the strong duality holds by Theorem 2. Thus, $\Theta(c_{*})=f_{*}$, which by [14, Corollary 3.3 and Remark 8] means that the penalty function $F$ is globally exact and $c_{*}\geq c_{*}(F)$. ∎
Remark 4.
(i) One can readily verify that in the case of penalty function (13) the lower semicontinuity of the perturbation function $\beta$ is equivalent to the strong duality (see, e.g. [15, Theorem 3.16]).
(ii) By the theorem above the global exactness of the penalty function $F$ is a necessary condition for the sequence of penalty parameters $\{c_{n}\}$ generated by the primal-dual penalty method to be bounded (or, equivalently, to converge). One can show that under some additional assumptions (in particular, the assumption that the sequence $\{\delta_{n}\}$ has a subsequence that is bounded away from zero) the global exactness of the penalty function $F$ is also sufficient for the boundedness of the sequence of penalty parameters. See [16, Theorem 5 and Remark 7] for more details.
4.2 Primal-dual rounded weighted $\ell_{1}$-penalty method
Let us now consider the primal-dual rounded weighted $\ell_{1}$-penalty method from [9] that can be viewed as a primal-dual method for the $\ell_{1}$-penalty function with multidimensional penalty parameter (i.e. individual penalty parameter for each constraint; see [16]) combined with smoothing techniques [29, 40, 27, 23, 20].
In this section we study a primal-dual method for the following mathematical programming problem:
$\min\>f(x)\quad\text{subject to}\quad h(x)=0,\quad g(x)\leq 0,\quad x\in Q,$ (17)
where $h\colon X\to\mathbb{R}^{m}$, $h=(h_{1},\ldots,h_{m})$, and $g\colon X\to\mathbb{R}^{\ell}$, $g=(g_{1},\ldots,g_{\ell})$, are given functions, and the inequality $g(x)\leq 0$ is understood coordinate-wise. This problem can be rewritten as the problem $(\mathcal{P})$, if one defines $Y=\mathbb{R}^{m+\ell}$ and $\Phi(\cdot):=\{h(\cdot)\}\times\{g(\cdot)+\mathbb{R}_{+}^{\ell}\}$. If the space $Y$ is equipped with the $\ell_{1}$-norm, then the infeasibility measure takes the following form:
$\operatorname{dist}(0,\Phi(x))=\sum_{i=1}^{m}|h_{i}(x)|+\sum_{j=1}^{\ell}\max\{0,g_{j}(x)\},\quad x\in X.$ (18)
The classic $\ell_{1}$-penalty function for problem (17) is defined as follows:
$F_{\ell_{1}}(x,c)=f(x)+c\Big{(}\sum_{i=1}^{m}|h_{i}(x)|+\sum_{j=1}^{\ell}\max\{0,g_{j}(x)\}\Big{)},$
where $c>0$ is the penalty parameter.
Burachik et al. in [9] proposed to consider a rounded weighted $\ell_{1}$-penalty function for problem (17), that is, they proposed to replace the nonsmooth functions $|h_{i}(\cdot)|$ and $\max\{0,g_{j}(\cdot)\}$ with their smoothing approximations (see [29, 40, 27, 23, 20] for more details on smoothing approximations of nonsmooth penalty functions) and to use individual penalty parameter for each of the constraint $h_{i}(x)=0$ and $g_{j}(x)\leq 0$ instead of the single penalty parameter $c$. To define this penalty function, first for any $w>0$ we introduce the following smoothing approximations of the functions $|\cdot|$ and $\max\{0,\cdot\}$:
$\eta(t,w)=\begin{cases}\frac{t^{2}}{2w},&\text{if }|t|<w,\\ |t|-\frac{w}{2},&\text{if }|t|\geq w,\end{cases}\qquad\gamma(t,w)=\begin{cases}\frac{t^{2}}{2w},&\text{if }0<t<w,\\ t-\frac{w}{2},&\text{if }t\geq w,\\ 0,&\text{if }t\leq 0.\end{cases}$
By definition we set $\eta(t,0)=|t|$ and $\gamma(t,0)=\max\{0,t\}$ for all $t\in\mathbb{R}$.
Introduce the rounded weighted $\ell_{1}$-penalty function for problem (17) as follows:
$F\big{(}x,(u,v,w)\big{)}:=f(x)+\sum_{i=1}^{m}u_{i}\eta(h_{i}(x),w)+\sum_{j=1}^{\ell}v_{j}\gamma(g_{j}(x),w)$
Here $u\in\mathbb{R}^{m}_{+}$ and $v\in\mathbb{R}^{\ell}_{+}$ are multidimensional penalty parameters and $w\geq 0$ is the smoothing parameter. In this case one can define $M=\mathbb{R}^{m+\ell+1}$,
$\varphi(x,\mu)=\sum_{i=1}^{m}u_{i}\eta(h_{i}(x),w)+\sum_{j=1}^{\ell}v_{j}\gamma(g_{j}(x),w),$ (19)
for any $x\in X$ and $\mu=(u,v,w)\in M$, and the dual function
$\Theta(\mu)=\inf_{x\in Q}\Big{(}f(x)+\sum_{i=1}^{m}u_{i}\eta(h_{i}(x),w)+\sum_{j=1}^{\ell}v_{j}\gamma(g_{j}(x),w)\Big{)}.$
The function $\Theta$ is obviously usc as the infimum of a family of continuous functions, while the function $(u,v)\mapsto\Theta(u,v,w)$ is concave as the infimum of a family of affine functions. Arguing in the same way as in the previous section one can verify that if $x_{n}$ is an $\varepsilon_{n}$-optimal solution of the problem
$\min_{x}\enspace F(x,\mu_{n})\quad\text{subject to}\quad x\in Q$ (20)
for some $\varepsilon_{n}\geq 0$ and $\mu_{n}=(u_{n},v_{n},w_{n})\in M$, then the vector $p_{n}\in\mathbb{R}^{m+\ell}$,
$p_{n}=\Big{(}\eta(h_{1}(x_{n}),w_{n}),\ldots,\eta(h_{m}(x_{n}),w_{n}),\gamma(g_{1}(x_{n}),w_{n}),\ldots,\gamma(g_{\ell}(x_{n}),w_{n})\Big{)}$
is an $\varepsilon_{n}$-supergradient of the concave function $(u,v)\mapsto\Theta(u,v,w_{n})$ (note that each coordinate of the vector $p_{n}$ is finite, since the optimal value of problem (20) does not exceed $f_{*}$). Therefore one can use the following subgradient-type step as an update of the multidimensional penalty parameter:
$(u_{n+1},v_{n+1})=(u_{n},v_{n})+s_{n}\big{(}p_{n}+\delta_{n}\big{)},$
where $s_{n}>0$ is the step size, while $\delta_{n}\geq 0$ is a correction. Utilising this penalty parameter update and taking into account the fact that $\eta(\cdot,w)\to|\cdot|$ and $\gamma(\cdot,w)\to\max\{0,\cdot\}$ as $w\to 0$ uniformly on $\mathbb{R}$, one can define the following primal-dual rounded weighted $\ell_{1}$-penalty method given in Algorithm 3 that was first introduced by Burachik et al. in [9] in the case when $\varepsilon_{n}\equiv 0$ and the set $Q$ is compact.
Initialization. Choose an initial value of the multidimensional penalty parameters $u_{0}\geq 0$ and $v_{0}\geq 0$, an initial value of the smoothing parameter $w_{0}>0$, and a bounded above sequence of tolerances $\{\varepsilon_{n}\}\subset[0,+\infty)$. Put $n=0$.
Step 1. Solution of subproblem. Find an $\varepsilon_{n}$-optimal solution $x_{n}$ of the problem
$\min_{x}\>f(x)+\sum_{i=1}^{m}u_{ni}\eta(h_{i}(x),w_{n})+\sum_{j=1}^{\ell}v_{nj}\gamma(g_{j}(x),w_{n})\quad\text{s.t.}\quad x\in Q,$
where $u_{n}=(u_{n1},\ldots,u_{nm})$ and $v_{n}=(v_{n1},\ldots,v_{n\ell})$.
Step 2. Update of parameters. Compute the vector
$p_{n}=\Big{(}\eta(h_{1}(x_{n}),w_{n}),\ldots,\eta(h_{m}(x_{n}),w_{n}),\gamma(g_{1}(x_{n}),w_{n}),\ldots,\gamma(g_{\ell}(x_{n}),w_{n})\Big{)},$
choose some $s_{n}>0$, $\delta_{n}\in\mathbb{R}^{m+\ell}_{+}$, and define
$(u_{n+1},v_{n+1})=(u_{n},v_{n})+s_{n}\big{(}p_{n}+\delta_{n}\big{)}.$
Choose some $w_{n+1}\leq w_{n}$, increment $n$, and go to Step 1.
Algorithm 3 Primal-dual rounded weighted-$\ell_{1}$ penalty method
Let us show how one can easily prove convergence of the primal-dual rounded weighted $\ell_{1}$-penalty method with the use of the universal convergence theorem for inexact primal-dual methods. For the sake of convenience, denote $\mu_{n}=(u_{n},v_{n},w_{n})$.
Theorem 4.
Let the penalty function $F(\cdot,\mu_{0})$ be bounded below and the sequence $\{(x_{n},u_{n},v_{n},w_{n})\}$ be generated by the primal-dual rounded weighted $\ell_{1}$-penalty method. Suppose also that one of the following two conditions is valid:
  1. 1.
    there exists $s_{*}>0$ such that $s_{n}\geq s_{*}$ for all $n\in\mathbb{N}$;
  2. 2.
    there exists a nondecreasing function $\omega\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $\omega(t)>0$ for any $t>0$, and $s_{n}\geq\omega(\|p_{n}\|)$ for all $n\in\mathbb{N}$, where $\|\cdot\|$ is the Euclidean norm.
Then $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$ and
$\vartheta_{*}\leq\liminf_{n\to\infty}f(x_{n})$ $\leq\limsup_{n\to\infty}f(x_{n})\leq\vartheta_{*}+\varepsilon_{*},$
$\vartheta_{*}-\varepsilon_{*}\leq\liminf_{n\to\infty}\Theta(\mu_{n})$ $\leq\limsup_{n\to\infty}\Theta(\mu_{n})\leq\vartheta_{*},$
where $\vartheta_{*}=\min\{f_{*},\liminf_{y\to 0}\beta(y)\}$. In addition, if the optimal value function $\beta$ be lsc at the origin, then
$\limsup_{n\to\infty}\Big{(}\sum_{i=1}^{m}u_{ni}\eta(h_{i}(x_{n}),w_{n})+\sum_{j=1}^{\ell}v_{nj}\gamma(g_{j}(x_{n}),w_{n})\Big{)}\leq\varepsilon_{*}$ (21)
and the following statements hold true:
  1. 1.
    if the function $h$ is continuous on $Q$, and the functions $f$ and $g_{j}$, $j\in\{1,\ldots,\ell\}$, are lsc on $Q$, then any limit point of the sequence $\{x_{n}\}$ (if such point exists) is an $\varepsilon_{*}$-optimal solution of problem (17);
  2. 2.
    if the sequence $\{(u_{n},v_{n})\}$ is bounded, then the sequence $\{(u_{n},v_{n},w_{n})\}$ converges and its limit point is an $\varepsilon_{*}$-optimal solutions of the dual problem;
  3. 3.
    in the case $\varepsilon_{*}=0$ for the boundedness of the sequence $\{(u_{n},v_{n})\}$ it is necessary that the $\ell_{1}$-penalty function for problem (17) is globally exact.
Proof.
Note that $u_{n}\leq u_{n+1}$ and $v_{n}\leq v_{n+1}$ for any $n\in\mathbb{N}$, since the functions $\eta$ and $\gamma$ are nonnegative. Furthermore, the functions $\eta(t,w)$ and $\gamma(t,w)$ are nonincreasing in $w$ (see [9]). Hence taking into account the way in which the multidimensional penalty parameters $u_{n}$ and $v_{n}$ and the smoothing parameter $w_{n}$ are updated by the method, one can conclude that $F(\cdot,\mu_{n})\leq F(\cdot,\mu_{n+1})$ for any $n\in\mathbb{N}$. Therefore, the boundedness below of the function $F(\cdot,\mu_{0})$ guarantees that Assumption 1 holds true.
The function $\varphi$ defined in (19) is obviously nonnegative. Moreover, for any $\mu\in M$ and any feasible point $x$ one has $\varphi(x,\mu)=0$ . Consequently, by Lemma 1 the weak duality holds and $\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0$. Moreover, the sequence $\{\mu_{n}\}$ clearly satisfies the $\varphi$-convergence to zero condition with any sequence $\{t_{n}\}\subset(0,+\infty)$ such that
$\lim_{n\to\infty}\Big{(}\sum_{i=1}^{m}u_{ni}\eta(t_{n},w_{n})+\sum_{j=1}^{\ell}v_{nj}\gamma(t_{n},w_{n})\Big{)}=0$
(at least one such sequence $\{t_{n}\}$ obviously exists). Thus, in order to apply the universal convergence theorem for the inexact primal-dual method it remains to check that $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$.
Suppose by contradiction that the sequence $\{\operatorname{dist}(0,\Phi(x_{n}))\}$ does not converge to zero. Then taking into account (18) one can conclude that either the sequence $\{|h_{r}(x_{n})|\}$ does not converge to zero for some $r\in\{1,\ldots,m\}$ or the sequence $\{\max\{0,g_{r}(x_{n})\}\}$ does not converge to zero for some $r\in\{1,\ldots,\ell\}$. We will consider only the first case, since the proof in the second case almost literally repeats the proof in the first one.
Thus, we assume that the sequence $\{|h_{r}(x_{n})|\}$ does not converge to zero for some $r\in\{1,\ldots,m\}$. Then the sequence $\{\eta(h_{r}(x_{n}),w_{0})\}$ does not converge to zero either and
$\sum_{n=0}^{\infty}\eta(h_{r}(x_{n}),w_{0})=+\infty.$ (22)
According to Step 2 of Algorithm 3 one has
$u_{nr}=u_{0r}+\sum_{k=0}^{n-1}s_{k}\big{(}\eta(h_{r}(x_{k}),w_{k})+\delta_{kr}\big{)}\geq u_{0r}+\sum_{k=0}^{n-1}s_{k}\eta(h_{r}(x_{k}),w_{0})$
(here we used the facts that $\delta_{n}\geq 0$, $w_{n+1}\leq w_{n}$ for all $n$, and the function $\eta(t,w)$ is nonincreasing in $w$). If $s_{n}\geq s_{*}>0$ for all $n\in\mathbb{N}$, then $u_{nr}\to+\infty$ as $n\to+\infty$ due to (22). In turn, if $s_{n}\geq\omega(\|p_{n}\|)$ for all $n\in\mathbb{N}$, then $s_{n}\geq\omega(\eta(h_{r}(x_{n}),w_{0}))$ due to the fact that the function $\omega$ is nondecreasing. Clearly, the sequence $\{\omega(\eta(h_{r}(x_{n}),w_{0}))\eta(h_{r}(x_{k}),w_{0})\}$ does not converge to zero, since the sequence $\{\eta(h_{r}(x_{k}),w_{0})\}$ does not converge to zero, the function $\omega$ is nondecreasing, and $\omega(t)>0$ for any $t>0$. Therefore,
$\lim_{n\to\infty}u_{nr}$ $\geq u_{0r}+\lim_{n\to\infty}\sum_{k=0}^{n-1}\omega(\eta(h_{r}(x_{n}),w_{0}))\eta(h_{r}(x_{k}),w_{0})$
$=u_{0r}+\sum_{k=0}^{\infty}\omega(\eta(h_{r}(x_{n}),w_{0}))\eta(h_{r}(x_{k}),w_{0})=+\infty.$
Thus, in both cases $u_{nr}\to+\infty$ as $n\to\infty$.
Bearing in mind the facts that $u_{n}\geq u_{0}$, $v_{n}\geq v_{0}$ and $w_{n}\leq w_{0}$, and the functions $\eta(t,w)$ and $\gamma(t,w)$ are nonincreasing in $w$ one obtains:
$F(x_{n},\mu_{n})$ $=F(x_{n},\mu_{0})\begin{aligned} &+\sum_{i=1}^{m}\big{(}u_{ni}\eta(h_{i}(x_{n}),w_{n})-u_{0i}\eta(h_{i}(x_{n}),w_{0})\big{)}\\ &+\sum_{j=1}^{\ell}\big{(}v_{nj}\gamma(g_{j}(x_{n}),w_{n})-v_{0j}\eta(g_{j}(x_{n}),w_{0})\big{)}\end{aligned}$
$\geq\inf_{x\in Q}F(x,\mu_{0})+\big{(}u_{nr}-u_{0r}\big{)}\eta(h_{r}(x_{n}),w_{0}).$
Therefore
$\limsup_{n\to\infty}F(x_{n},\mu_{n})=+\infty$ (23)
due to the facts that $u_{nr}\to+\infty$ as $n\to\infty$ and the sequence $\{\eta(h_{r}(x_{n}),w_{0})\}$ does not converge to zero. However, by the definition of $x_{n}$ and the fact that $F(x,\mu_{n})=f(x)$ for any feasible $x$ one has $F(x_{n},\mu_{n})\leq f_{*}+\varepsilon_{n}$ for all $n\in\mathbb{N}$, which contradicts (23).
Thus, $\operatorname{dist}(0,\Phi(x_{n}))\to 0$ as $n\to\infty$ and all claims of the theorem, except for the claim about the case $\varepsilon_{*}=0$, hold true by Theorem 1 (one only needs to note that the convergence of the sequence $\{(u_{n},v_{n},w_{n})\}$ is equivalent to the boundedness of the sequence $\{(u_{n},v_{n})\}$ due to the facts that $0\leq w_{n+1}\leq w_{n}$, $u_{n}\leq u_{n+1}$, and $v_{n}\leq v_{n+1}$ for all $n\in\mathbb{N}$).
Let us finally show that in the case $\varepsilon_{*}=0$ for the boundedness of the sequence $\{(u_{n},v_{n})\}$ it is necessary that the $\ell_{1}$-penalty function for problem (17) is globally exact. Indeed, if this sequence is bounded, then the sequence $\{(u_{n},v_{n},w_{n})\}$ converges to some point $\mu_{*}=(u_{*},v_{*},w_{*})$. By Theorem 2 this point is a globally optimal solution of the dual problem and the strong duality holds true, which implies that $\Theta(\mu_{*})=f_{*}$.
Let $c_{*}=\max\{(u_{*})_{1},\ldots,(u_{*})_{m},(v_{*})_{1},\ldots,(v_{*})_{\ell}\}$. Then
$F_{\ell_{1}}(x,c_{*})\geq F(x,\mu_{*})\geq\Theta(\mu_{*})=f_{*}\quad\forall x\in Q$
(the first inequality follows from the properties of the functions $\eta$ and $\gamma$; see [9]). By [14, Corollary 3.3 and Remark 8] the inequality above implies that the penalty function $F_{\ell_{1}}$ is globally exact. Thus, the boundedness of the sequence $\{(u_{n},v_{n})\}$ implies that the $\ell_{1}$-penalty function for problem (17) is globally exact. ∎
Let us also show how the previous theorem can be strengthened in the case, when $Q$ is a compact set and the method is asymptotically exact, that was considered in [9].
Corollary 4.
Let the set $Q$ be compact, the function $h$ be continuous on $Q$, the functions $f$ and $g_{j}$, $j\in\{1,\ldots,\ell\}$, be lsc on $Q$, and the sequence $\{(x_{n},u_{n},v_{n},w_{n})\}$ be generated by the primal-dual rounded weighted $\ell_{1}$-penalty method with $\varepsilon_{n}\to 0$ as $n\to\infty$. Suppose also that one of the following two conditions is valid:
  1. 1.
    there exists $s_{*}>0$ such that $s_{n}\geq s_{*}$ for all $n\in\mathbb{N}$;
  2. 2.
    there exists a nondecreasing function $\omega\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $\omega(t)>0$ for any $t>0$, and $s_{n}\geq\omega(\|p_{n}\|)$ for all $n\in\mathbb{N}$, where $\|\cdot\|$ is the Euclidean norm.
Then
$\lim_{n\to\infty}\operatorname{dist}(0,\Phi(x_{n}))=0,\quad\lim_{n\to\infty}f(x_{n})=\lim_{n\to\infty}\Theta(\mu_{n})=f_{*},$
the sequence $\{x_{n}\}$ is bounded, and all its limit points are globally optimal solutions of the problem $(\mathcal{P})$. If, in addition, the sequence $\{(u_{n},v_{n})\}$ is bounded, then the sequence $\{(u_{n},v_{n},w_{n})\}$ converges to a globally optimal solution of the dual problem.
Proof.
Under the assumptions of the corollary the function $F(\cdot,\mu_{0})$ is obviously bounded below on $Q$. Let us check that the optimal value function $\beta$ is lsc at the origin. Then applying Theorem 4 with $\varepsilon_{*}=0$ we arrive at the desired result.
Let $\{y_{n}\}\subset Y:=\mathbb{R}^{m+\ell}$ be any sequence such that $y_{n}\to 0$ as $n\to\infty$ and
$\beta_{*}:=\lim_{n\to\infty}\beta(y_{n})=\liminf_{y\to 0}\beta(y).$
Note that the sequence $\{\beta(y_{n})\}$ is bounded below due to the compactness of the set $Q$. If there does not exist $n_{0}\in\mathbb{N}$ such that the sequence $\{\beta(y_{n})\}_{n\geq n_{0}}$ is bounded above, then $\beta_{*}=+\infty$ and the function $\beta$ is obviously lsc at the origin. Therefore, without loss of generality one can suppose that the sequence $\{\beta(y_{n})\}$ is bounded.
By the definition of $\beta$ for any $n\in\mathbb{N}$ one can find $x_{n}\in Q$ such that
$\beta(y_{n})\leq f(x_{n})+\varepsilon_{n},\quad$ $h_{i}(x_{n})=y_{ni},\quad i\in\{1,\ldots,m\},$
$g_{j}(x_{n})\leq y_{n(m+j)},\quad j\in\{1,\ldots,\ell\}.$
Since $Q$ is compact, one can find a subsequence $\{x_{n_{k}}\}$ that converges to some $x_{*}\in Q$. From the inequalities above, the assumptions on the function $h$ and $g$, and the fact that $y_{n}\to 0$ as $n\to\infty$ it follows that $h(x_{*})=0$ and $g(x_{*})\leq 0$, that is, the point $x_{*}$ is feasible for problem (17). Moreover, due to the lower semicontinuity of $f$ one has
$\liminf_{y\to 0}\beta(y)=\lim_{n\to\infty}\beta(y_{n})=\lim_{n\to\infty}f(x_{n})\geq f(x_{*})\geq f_{*}=:\beta(0),$
which means that the optimal value function $\beta$ is lsc at the origin. ∎
Remark 5.
Let us note that [9, Theorem 4.1] is a particular case of the corollary above. Furthermore, this corollary and Theorem 4 significantly generalize the convergence analysis presented in [9], since we do not assume that the set $Q$ is compact, $\varepsilon_{n}\equiv 0$, and either $s_{n}\equiv 1/\|p_{n}\|^{2}$ or $s_{n}\equiv\|p_{n}\|^{2}$ (in the first case $s_{n}\geq s_{*}>0$ for all $n\in\mathbb{N}$, while in the second case $s_{n}\geq\omega(\|p_{n}\|)$ for $\omega(t)=t^{2}$). Thus, with the use of the universal convergence theorem for inexact primal-dual methods we not only managed to recover existing results for the primal-dual rounded weighted $\ell_{1}$-penalty method, but also significantly generalized these results.
4.3 Augmented Lagrangian method for cone constrained optimization
In this section we study the classic augmented Lagrangian method based on the Hestenes-Powell-Rockafellar augmented Lagrangian [3]. This method has been extended, in particular, to the case of optimization problems with nonlinear second order cone constraints [22] and nonlinear semidefinite programming problems [26]. Our aim is to extend this method to the case of general cone constrained optimization problems (which to the best of the author’s knowledge has never been done before) and prove its global convergence with the use of the universal convergence theorem for inexact primal-dual methods.
Let $Y$ be a Hilbert space with inner product $\langle\cdot,\cdot\rangle$, $K\subset Y$ be a closed convex cone, and $G\colon X\to Y$ be a given function. In this section we consider the following cone constrained optimization problem:
$\min\>f(x)\quad\text{subject to}\quad G(x)\in K,\quad x\in Q.$ (24)
This problem can be rewritten as the problem $(\mathcal{P})$ with $\Phi(\cdot):=G(\cdot)-K$. Note that the infeasibility measure satisfies the equality $\operatorname{dist}(0,\Phi(x))=\operatorname{dist}(G(x),K)$.
Denote by $K^{*}=\{y^{*}\in Y\mid\langle y^{*},y\rangle\leq 0\>\forall y\in K\}$ the polar cone of the cone $K$. Since $Y$ is a Hilbert space, for any closed convex set $V\subseteq Y$ and any $y\in Y$ there exists a unique metric projection of $y$ onto $V$, which we denote by $\Pi_{V}(y)$. Following [32, Section 11.K] and [34] one can define an augmented Lagrangian for problem (24) as follows:
$\begin{split}\mathscr{L}(x,\lambda,c)&=f(x)+\frac{1}{2c}\Big{[}\big{\|}\Pi_{K^{*}}(\lambda+cG(x))\big{\|}^{2}-\|\lambda\|^{2}\Big{]}\\ &=f(x)+\frac{1}{2c}\Big{[}\operatorname{dist}\big{(}\lambda+cG(x),K\big{)}^{2}-\|\lambda\|^{2}\Big{]},\end{split}$ (25)
It can be viewed an extension of the Hestenes-Powell-Rockafellar augmented Lagrangian [31, 3] to the case of general cone constrained optimization problems. Here $\lambda\in Y$ is a multiplier and $c>0$ is a penalty parameter. Note that the dual function
$\Theta(\lambda,c)=\inf_{x\in Q}\mathscr{L}(x,\lambda,c)$
is concave in $(\lambda,c)$ and usc (see [32, Section 11.K] and [34]).
Let us now describe a basic augmented Lagrangian method. Its theoretical scheme is given in Algorithm 4.
Initialization. Choose initial values of the multiplier $\lambda_{0}\in K^{*}$ and the penalty parameter $c_{0}>0$, and a bounded above sequence of tolerances $\{\varepsilon_{n}\}\subset[0,+\infty)$. Put $n=0$.
Step 1. Solution of subproblem. Find an $\varepsilon_{n}$-optimal solution $x_{n}$ of the problem
$\min_{x}\>\mathscr{L}(x,\lambda_{n},c_{n})\quad\text{subject to}\quad x\in Q.$
Step 2. Multiplier update. Choose some $\lambda_{n+1}\in K^{*}$.
Step 3. Penalty parameter update. Choose some $c_{n+1}\geq c_{n}$, increment $n$, and go to Step 1.
Algorithm 4 Inexact basic augmented Lagrangian method
In order to incorporate several particular cases into one method, we do not specify the way in which the multipliers and the penalty parameter are updated in this method. In particular, one can define
$\lambda_{n+1}=\lambda_{n}+\Big{(}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\Big{)}$
as in [3, Algorithm 4.1], [22], [24, Algorithm 1], [26, Algorithm 1], or one can put
$\lambda_{n+1}=\Pi_{T\cap K^{*}}\left[\lambda_{n}+\Big{(}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\Big{)}\right]$ (26)
for some convex bounded set $T\subset Y$ containing a neighbourhood of zero, as in the multiplier method with safeguarding (see [24, Algorithm 2] and [26, Algorithm 2]).
The penalty parameter can be updated according to the following standard rule (see [3, 24, 26]):
$c_{n+1}=\begin{cases}c_{n},&\text{if }\sigma_{n}\leq\tau\sigma_{n-1},\\ \gamma c_{n},&\text{otherwise}\end{cases}\quad\forall n\in\mathbb{N},$ (27)
for some fixed $\tau\in(0,1)$ and $\gamma>1$, where
$\sigma_{n}:=\frac{1}{c_{n}}\big{\|}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\big{\|}$
for any $n\in\mathbb{N}$ and $\sigma_{-1}:=+\infty$. Note that if rule (27) is used, then $\sigma_{n}\to 0$ as $n\to\infty$, if the sequence $\{c_{n}\}$ is bounded.
Let us prove convergence of the inexact basic augmented Lagrangian method with the use of the universal convergence theorem. For the sake of convenience, denote
$\begin{split}\varphi(x,(\lambda,c))&=\Big{(}\frac{1}{2c}\Big{[}\big{\|}\Pi_{K^{*}}(\lambda+cG(x))\big{\|}^{2}-\|\lambda\|^{2}\Big{]}\\ &=\frac{1}{2c}\Big{[}\operatorname{dist}\big{(}\lambda+cG(x),K\big{)}^{2}-\|\lambda\|^{2}\Big{]}\end{split}$ (28)
and $\mu_{n}=(\lambda_{n},c_{n})$.
Theorem 5.
Let $\{(x_{n},\lambda_{n},c_{n})\}$ be the sequence generated by the inexact basic augmented Lagrangian method. Suppose that the sequence $\{f(x_{n})\}$ is bounded below and the two following assumptions hold true:
  1. 1.
    if $c_{n}\to+\infty$ as $n\to\infty$, then
    $\lim_{n\to\infty}\frac{\|\lambda_{n}\|}{\sqrt{c_{n}}}=0.$ (29)
  2. 2.
    if the sequence $\{c_{n}\}$ is bounded, then (a) the sequence $\{\lambda_{n}\}$ is bounded, (b) the sequence $\sigma_{n}:=(1/c_{n})\|\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\|$, $n\in\mathbb{N}$, converges to zero, and (c) either the sequence $\{\Pi_{K}(G(x_{n}))\}$ is bounded or $Y=\mathbb{R}^{m+\ell}$ and $K=\{0\}\times\mathbb{R}^{\ell}_{-}$, where $\mathbb{R}_{-}=(-\infty,0]$.
Then
$\lim_{n\to\infty}\operatorname{dist}(G(x_{n}),K)=0,\quad\limsup_{n\to\infty}\varphi(x_{n},(\lambda_{n},c_{n}))\leq\varepsilon_{*},$
$\vartheta_{*}\leq\liminf_{n\to\infty}f(x_{n})\leq\limsup_{n\to\infty}f(x_{n})\leq\vartheta_{*}+\varepsilon_{*},$
$\vartheta_{*}-\varepsilon_{*}\leq\liminf_{n\to\infty}\Theta(\lambda_{n},c_{n})\leq\limsup_{n\to\infty}\Theta(\lambda_{n},c_{n})\leq\vartheta_{*},$
where $\vartheta_{*}:=\min\{f_{*},\liminf_{y\to 0}\beta(y)\}$. If, in addition, the optimal value function $\beta$ is lsc at the origin, then $\vartheta_{*}=f_{*}$ and the following statements hold true:
  1. 1.
    if $f$ is lsc on $Q$ and $G$ is continuous on $Q$, then any limit point of the sequence $\{x_{n}\}$ (if such point exists) is an $\varepsilon_{*}$-optimal solution of problem (24);
  2. 2.
    if the sequence of penalty parameters $\{c_{n}\}$ is bounded, then any limit point of the sequence $\{(\lambda_{n},c_{n})\}$ (if such point exists) is an $\varepsilon_{*}$-optimal solution of the dual problem.
Proof.
To apply the universal convergence theorem, define $M=Y\times(0,+\infty)$ and $F(x,\mu)=f(x)+\varphi(x,\mu)$ with $\mu=(\lambda,c)$. Our aim is to show that
$\lim_{n\to\infty}\operatorname{dist}(0,\Phi(x_{n}))=0,\quad\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0,$ (30)
the weak duality holds true, and the sequence $\{(\lambda_{n},c_{n})\}$ satisfies the $\varphi$-convergence to zero condition. Then applying the universal convergence theorem for inexact primal-dual methods one obtains the required result.
Part 1. Observe that for any feasible point $x$ and any $\lambda\in Y$ and $c>0$ one has
$\operatorname{dist}(\lambda+cG(x),K)=\inf_{y\in K}\|\lambda+cG(x)-y\|\leq\|\lambda+cG(x)-cG(x)\|\leq\|\lambda\|$
since $G(x)\in K$ and $K$ is a cone. Therefore, for any feasible point $x$ and any $\mu=(\lambda,c)\in M$ one has
$\varphi(x,\mu)=\frac{1}{2c}\Big{[}\operatorname{dist}\big{(}\lambda+cG(x),K\big{)}^{2}-\|\lambda\|^{2}\Big{]}\leq\frac{1}{2c}\Big{[}\|\lambda\|^{2}-\|\lambda\|^{2}\Big{]}\leq 0$
(see (28)), which by Lemma 1 implies that the weak duality holds true.
Part 2. Let us now check the $\varphi$-convergence to zero condition. Indeed, fix some $n\in\mathbb{N}$, $t_{n}>0$, and $z_{n}\in Q$ such that $\operatorname{dist}(G(z_{n}),K)\leq t_{n}$. Then $\operatorname{dist}(c_{n}G(z_{n}),K)\leq c_{n}t_{n}$ and, therefore, one can find a vector $y_{n}\in K$ such that $\|c_{n}G(z_{n})-y_{n}\|\leq 2c_{n}t_{n}$. Consequently, one has
$\operatorname{dist}(\lambda_{n}+c_{n}G(z_{n}),K)\leq\|\lambda_{n}+c_{n}G(z_{n})-y_{n}\|\leq\|\lambda_{n}\|+2c_{n}t_{n},$
which implies that
$\operatorname{dist}(\lambda_{n}+c_{n}G(z_{n}),K)^{2}-\|\lambda_{n}\|^{2}$ $\leq\big{(}\|\lambda_{n}\|+2c_{n}t_{n}\big{)}^{2}-\|\lambda_{n}\|^{2}$
$=4c_{n}t_{n}\|\lambda_{n}\|+4c_{n}^{2}t_{n}^{2}.$
Therefore, choosing $t_{n}>0$ in such a way that $2t_{n}\|\lambda_{n}\|+2c_{n}t_{n}^{2}\leq 1/(n+1)$ one gets that for any sequence $\{z_{n}\}\subset Q$ with $\operatorname{dist}(G(z_{n}),K)\leq t_{n}$ one has
$\varphi(z_{n},\mu_{n})\leq\frac{1}{n+1}\quad\forall n\in\mathbb{N}$
(see (28)). Therefore, the sequence $\{(\lambda_{n},c_{n})\}$ satisfied the $\varphi$-convergence to zero condition for any such sequence $\{t_{n}\}$.
Part 3. Thus, it remains to check that conditions (30) hold true. Let us consider two cases corresponding to the two assumptions of the theorem.
Case I. Let the sequence of penalty parameters $\{c_{n}\}$ be bounded. Suppose that $\operatorname{dist}(G(x_{n}),K)\to 0$ as $n\to\infty$. From (28) it follows that
$\varphi(x_{n},\mu_{n})=\frac{1}{2c_{n}}\Big{(}\big{\|}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))\big{\|}-\|\lambda_{n}\|\Big{)}\Big{(}\operatorname{dist}\big{(}\lambda_{n}+c_{n}G(x_{n}),K\big{)}+\|\lambda_{n}\|\Big{)}$
The sequence $\{\operatorname{dist}(\lambda_{n}+c_{n}G(x_{n}),K)+\|\lambda_{n}\|\}$ is bounded due to the boundedness of the sequences $\{\lambda_{n}\}$ and $\{c_{n}\}$ and the fact that $\operatorname{dist}(G(x_{n}),K)\to 0$ as $n\to\infty$. Hence taking into account the facts that
$\frac{1}{2c_{n}}\Big{|}\big{\|}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))\big{\|}-\|\lambda_{n}\|\Big{|}\leq\frac{1}{2c_{n}}\Big{\|}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\Big{\|}=\frac{1}{2}\sigma_{n}$
and the sequence $\{\sigma_{n}\}$ converges to zero by our assumption, one can conclude that $\varphi(x_{n},\mu_{n})\to 0$ as $n\to\infty$. Therefore it remains to check that $\operatorname{dist}(G(x_{n}),K)\to 0$ as $n\to\infty$. Let us consider two subcases corresponding to the two different assumptions.
Case I.A. Suppose that the sequence $\{\Pi_{K}(G(x_{n}))\}$ is bounded. From the boundedness of the sequence $\{c_{n}\}$ and the convergence of the sequence $\{\sigma_{n}\}$ to zero it follows that
$\varkappa_{n}:=\|\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\|\xrightarrow[n\to\infty]{}0.$ (31)
By the definition of projection for any $n\in\mathbb{N}$ and $z\in K^{*}$ one has
$\|c_{n}G(x_{n})\|$ $=\big{\|}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\big{\|}$
$\leq\Big{\|}\lambda_{n}+c_{n}G(x_{n})-\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))\Big{\|}+\varkappa_{n}$
$\leq\big{\|}\lambda_{n}+c_{n}G(x_{n})-z\big{\|}+\varkappa_{n}.$
Since $K^{*}$ is a convex cone and $\{\lambda_{n}\}\subset K^{*}$, for any $n\in\mathbb{N}$ and $z\in K^{*}$ one has $\lambda_{n}+z\in K^{*}$. Therefore for any $n\in\mathbb{N}$ and $z\in K^{*}$ one has
$\|c_{n}G(x_{n})\|\leq\big{\|}\lambda_{n}+c_{n}G(x_{n})-(\lambda_{n}+z)\big{\|}+\varkappa_{n}=\big{\|}c_{n}G(x_{n})-z\big{\|}+\varkappa_{n},$
which yields
$\|G(x_{n})\|\leq\operatorname{dist}(G(x_{n}),K^{*})+\frac{\varkappa_{n}}{c_{n}}=\|\Pi_{K}(G(x_{n}))\|+\frac{\varkappa_{n}}{c_{n}}\quad\forall n\in\mathbb{N}.$
Hence by applying Moreau’s decomposition theorem for cones [28, 35] one obtains that
$\|\Pi_{K}(G(x_{n}))\|^{2}+\|\Pi_{K^{*}}(G(x_{n}))\|^{2}=\|G(x_{n})\|^{2}\\ \leq\|\Pi_{K}(G(x_{n}))\|^{2}+2\frac{\varkappa_{n}}{c_{n}}\|\Pi_{K}(G(x_{n}))\|+\frac{\varkappa_{n}^{2}}{c_{n}^{2}}$
for all $n\in\mathbb{N}$ or, equivalently,
$\|\Pi_{K^{*}}(G(x_{n}))\|^{2}\leq 2\frac{\varkappa_{n}}{c_{n}}\|\Pi_{K}(G(x_{n}))\|+\frac{\varkappa_{n}^{2}}{c_{n}^{2}},$
which with the use of (31) and the boundedness of the sequences $\{\Pi_{K}(G(x_{n}))\}$ and $\{c_{n}\}$ implies that $\operatorname{dist}(G(x_{n}),K)=\|\Pi_{K^{*}}(G(x_{n}))\|\to 0$ as $n\to\infty$.
Case I.B. Suppose that $Y=\mathbb{R}^{m+\ell}$ and $K=\{0\}\times\mathbb{R}^{\ell}_{-}$. Then $K^{*}=\mathbb{R}^{m}\times\mathbb{R}^{\ell}_{+}$ and for any $n\in\mathbb{N}$ one has
$\sigma_{n}=\frac{1}{c_{n}}\Big{\|}\Pi_{K^{*}}(\lambda_{n}+c_{n}G(x_{n}))-\lambda_{n}\Big{\|}=\\ \bigg{\|}\Big{(}G_{1}(x_{n}),\ldots,G_{m}(x_{n}),\max\left\{G_{m+1}(x_{n}),-\frac{1}{c_{n}}(\lambda_{n})_{m+1}\right\},\ldots,\\ \max\left\{G_{m+\ell}(x_{n}),-\frac{1}{c_{n}}(\lambda_{n})_{m+\ell}\right\}\bigg{\|}.$
Combining the equality above with the facts that $\sigma_{n}\to 0$ as $n\to\infty$ and $\lambda_{n}\in K^{*}$ one can readily verify that $\operatorname{dist}(G(x_{n}),K)\to 0$ as $n\to\infty$.
Case II. Suppose that $c_{n}\to+\infty$ as $n\to\infty$ and equality (29) holds true. Recall that
$\varphi\big{(}x,(\lambda,c)\big{)}=\inf_{y\in K-G(x)}\left(-\langle\lambda,y\rangle+\frac{c}{2}\|y\|^{2}\right)$
(see [32, Section 11.K] and [34]). For any $y\in Y$ one has
$\langle\lambda,y\rangle+\frac{c}{2}\|y\|^{2}\geq-\|\lambda\|\|y\|+\frac{c}{2}\|y\|^{2}\geq\inf_{t\geq 0}\left(-\|\lambda\|t+\frac{c}{2}t^{2}\right)=-\frac{\|\lambda\|^{2}}{2c}.$
Therefore
$\varphi(x_{n},\mu_{n})\geq-\frac{\|\lambda_{n}\|^{2}}{2c_{n}}\quad\forall n\in\mathbb{N}$
and $\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0$ due to equality (29).
Let us now show that $\operatorname{dist}(G(x_{n}),K)\to 0$ as $n\to\infty$. Suppose by contradiction that this statement is false. Then there exist $\delta>0$ and a subsequence $\{x_{n_{k}}\}$ such that $\operatorname{dist}(G(x_{n_{k}}),K)\geq\delta$ for all $k\in\mathbb{N}$. Taking into account (29) one can suppose that
$\operatorname{dist}\left(\frac{\lambda_{n_{k}}}{c_{n_{k}}}+G(x_{n_{k}}),K\right)\geq\frac{\delta}{2}\quad\forall k\in\mathbb{N}.$
Hence bearing in mind (25), for any $k\in\mathbb{N}$ one gets that
$\mathscr{L}(x_{n_{k}},\lambda_{n_{k}},c_{n_{k}})$ $=f(x_{n_{k}})+\frac{1}{2c_{n_{k}}}\Big{[}\operatorname{dist}\big{(}\lambda_{n_{k}}+c_{n_{k}}G(x_{n_{k}}),K\big{)}^{2}-\|\lambda_{n_{k}}\|^{2}\Big{]}$
$=f(x_{n_{k}})+\frac{c_{n_{k}}}{2}\operatorname{dist}\Big{(}\frac{\lambda_{n_{k}}}{c_{n_{k}}}+G(x_{n_{k}}),K\Big{)}^{2}-\frac{\|\lambda_{n_{k}}\|^{2}}{2c_{n_{k}}}$
$\geq f(x_{n_{k}})+\frac{c_{n_{k}}}{4}\delta-\frac{\|\lambda_{n_{k}}\|^{2}}{2c_{n_{k}}}.$
Consequently, $\mathscr{L}(x_{n_{k}},\lambda_{n_{k}},c_{n_{k}})\to+\infty$ as $k\to\infty$ due to equality (29) and the boundedness below of the sequence $\{f(x_{n_{k}})\}$, which contradicts the fact that $\mathscr{L}(x_{n_{k}},\lambda_{n_{k}},c_{n_{k}})\leq f_{*}+\varepsilon_{n_{k}}$ for all $k\in\mathbb{N}$ due to the weak duality. Therefore, $\operatorname{dist}(G(x_{n}),K)\to 0$ as $n\to\infty$.
Thus, in both cases conditions (30) are satisfied and applying Theorem 1 we arrive at the required results. ∎
Remark 6.
Let us briefly discuss the assumptions of Theorem 5:
  1. 1.
    In the case of inequality constrained problems, condition (29) always holds true, if one uses the so-called conditional multiplier updating (see e.g. [25, 24]). An extension of the conditional multplier updating to the case of general cone constrained problems is possible and is an interesting topic for future research.
  2. 2.
    The assumption on the boundedness of the sequence of multipliers $\{\lambda_{n}\}$ is common in the literature on augmented Lagrangian methods [1, 3, 25, 24, 26]. One can ensure the boundedness of this sequence by using the multiplier method with safeguarding (26).
  3. 3.
    The sequence $\{\sigma_{n}\}$ convergences to zero whenever the sequence of penalty parameters $\{c_{n}\}$ is bounded, if the standard rule (27) for updating the penalty parameter is used.
  4. 4.
    The assumption on the boundedness of the sequence $\{\Pi_{K}(G(x_{n}))\}$ is satisfied, provided the sequence $\{x_{n}\}$ is bounded and $G$ maps bounded sets into bounded sets. In particular, if $X$ is finite dimensional, it is sufficient to suppose that the sequence $\{x_{n}\}$ is bounded and the function $G$ is continuous.
  5. 5.
    Note that problem (24) with $Y=\mathbb{R}^{m+\ell}$ and $K=\{0\}\times\mathbb{R}^{\ell}_{-}$ is just a standard mathematical programming problems with $m$ equality and $\ell$ inequality constraints.
Remark 7.
Theorem 5 unifies, significantly generalizes, and strengthens many existing results on converges of primal-dual augmented Lagrangian methods based on the Hestenes-Powell-Rockafellar augmented Lagrangian. In particular, [3, Theorem 5.2], [24, Theorem 2.4 and 3.1] in the case when $\sigma(z)\equiv 0.5z^{2}$, and [26, Theorems 1, 3, and 4] follow directly from Theorem 5. Moreover, Theorem 5 strengthens the corresponding theorems from [24, 26], since we do not impose the assumption that the objective function $f$ is bounded below on $Q$ and only suppose that the sequence $\{f(x_{n})\}$ is bounded below.
4.4 Multiplier algorithm based on PALM
Let us finally apply the universal convergence theorem for inexact primal-dual methods to the multiplier algorithm based on the so-called P-type Augmented Lagrangian Method (PALM) that was proposed in [37]. One can prove convergence of all other augmented Lagrangian methods from [37] with the use of the universal convergence theorem in exactly the same way.
In this section we study an augmented Lagrangian method for the inequality constrained problem:
$\min\>f(x)\quad\text{subject to}\quad g(x)\leq 0,\quad x\in Q,$ (32)
where $g\colon X\to\mathbb{R}^{\ell}$ is a given function. Fix some $a\in(0,+\infty]$, and let a function $\colon(-\infty,a)\times\mathbb{R}_{+}\to\mathbb{R}, P=P$(s,t)P=P(s,t), be continuous in $(s,t)$ and continuously differentiable in $s$. Suppose also that for some continuous function $r\colon\mathbb{R}_{+}\to\mathbb{R}$ one has
$P(s,t)\geq r(t)\quad\forall s\in(-\infty,a),\>t\in\mathbb{R}_{+}.$ (33)
Following [37], we call the function
$L_{P}(x,\lambda,c)=\begin{cases}f(x)+\frac{1}{c}\sum_{i=1}^{\ell}P(cg_{i}(x),\lambda_{i}),&\text{if }cg_{i}(x)<a\enspace\forall i\in\{1,\ldots,\ell\},\\ +\infty,&\text{otherwise}\end{cases}$
the P-type augmented Lagrangian for problem (32). Here $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\in\mathbb{R}^{\ell}_{+}$ are multipliers and $c>0$ is the penalty parameter. Several particular examples of the function $P$ can be found in [37, Section 2]. Here we only note that the well-known Hestenes-Powell-Rockafellar augmented Lagrangian [31, 3] is a particular case of the P-type augmented Lagrangian.
In [37, Section 4] the following primal-dual method, given in Algorithm 5 and called the multiplier algorithm based on PALM, was proposed.
Initialization. Choose initial points $x_{0}\in Q$, $\lambda_{0}\geq 0$, and $c_{0}>0$, and a bounded above sequence of tolerances $\{\varepsilon_{n}\}\subset[0,+\infty)$. Put $n=0$.
Step 1. Update of parameters. Compute
$\lambda_{(n+1)i}$ $=P^{\prime}_{s}(c_{n}g_{i}(x_{n}),\lambda_{ni}),\quad i\in\{1,\ldots,\ell\},$
$c_{n+1}$ $\geq(n+1)\max\Big{\{}1,\sum_{i=1}^{\ell}|r(\lambda_{(n+1)i})|\Big{\}},$
where $P^{\prime}_{s}$ is the derivative of the function $s\mapsto P(s,t)$.
Step 2. Solution of subproblem. Find an $\varepsilon_{n+1}$-optimal solution $x_{n+1}$ of the problem
$\min_{x}\>L_{P}(x,\lambda_{n},c_{n})\quad\text{subject to}\quad x\in Q.$
Increment $n$ and go to Step 1.
Algorithm 5 Multiplier algorithm based on PALM
Let us analyse convergence of Algorithm 5 with the use of the universal convergence theorem for inexact primal-dual methods.
Theorem 6.
Let $\{(x_{n},\lambda_{n},c_{n})\}$ be the sequence generated by Algorithm 5. Suppose that the sequence $\{f(x_{n})\}$ is bounded below and the function $P$ satisfies the following assumptions:
  1. 1.
    the function $s\mapsto P(s,t)$ is nondecreasing and $P(0,t)=0$ for any $t\geq 0$;
  2. 2.
    if $a=+\infty$, then $P(s,t)/s\to+\infty$ as $s\to+\infty$ for any $t\geq 0$.
Then
$\lim_{n\to\infty}\sum_{i=1}^{\ell}\max\{0,g_{i}(x_{n})\}=0,$ (34)
$\vartheta_{*}\leq\liminf_{n\to\infty}f(x_{n})\leq\limsup_{n\to\infty}f(x_{n})\leq\vartheta_{*}+\varepsilon_{*},$ (35)
$\vartheta_{*}-\varepsilon_{*}\leq\liminf_{n\to\infty}\Theta(\mu_{n})\leq\limsup_{n\to\infty}\Theta(\mu_{n})\leq\vartheta_{*},$ (36)
where $\vartheta_{*}=\min\{f_{*},\liminf_{y\to 0}\beta(y)\}$. If, in addition, the optimal value function $\beta$ is lsc at the origin and the functions $f$ and $g_{i}$, $i\in\{1,\ldots,\ell\}$ are lsc on $Q$, then $\vartheta_{*}=f_{*}$ and limit points of the sequence $\{x_{n}\}$ (if such points exist) are $\varepsilon_{*}$-optimal solutions of problem (32).
Proof.
It should be noted that the first assumption on the function $P$ ensures that $\lambda_{n}\geq 0$ for any $n\in\mathbb{N}$ and, therefore, the method is correctly defined.
Denote $M=\mathbb{R}^{\ell}_{+}\times(0,+\infty)$ and for any $\mu=(\lambda,c)\in M$ define
$\varphi(x,\mu)=\begin{cases}\frac{1}{c}\sum_{i=1}^{\ell}P(cg_{i}(x),\lambda_{i}),&\text{if }cg_{i}(x)<a\enspace\forall i\in\{1,\ldots,\ell\},\\ +\infty,&\text{otherwise.}\end{cases}$
Then $L_{P}(x,\lambda,c)=f(x)+\varphi(x,(\lambda,c))$. For the sake of convenience, denote $\mu_{n}=(\lambda_{n},c_{n})$.
By the first assumption on the function $P$ for any $s\leq 0$ and $t\in\mathbb{R}_{+}$ one has $P(s,t)\leq 0$, which implies that for any feasible point $x$ and any $\mu\in M$ one has $\varphi(x,\mu)\leq 0$. Hence by Lemma 1 the weak duality holds and, therefore, $L_{P}(x_{n},\lambda_{n},c_{n})\leq f_{*}+\varepsilon_{n}$ for any $n\in\mathbb{N}$.
Bearing in mind condition (33) and the rule for updating the penalty parameter on Step 1 of Algorithm 5 one gets that
$\varphi(x_{n},\mu_{n})=\frac{1}{c_{n}}\sum_{i=1}^{\ell}P(c_{n}g_{i}(x_{n}),\lambda_{ni})\geq\frac{1}{c_{n}}\sum_{i=1}^{\ell}r(\lambda_{ni})\geq-\frac{1}{c_{n}}\sum_{i=1}^{\ell}|r(\lambda_{ni})|\geq-\frac{1}{n}$
for any $n\geq 1$, which implies that $\liminf_{n\to\infty}\varphi(x_{n},\mu_{n})\geq 0$.
Let us now check that $\max\{0,g_{i}(x_{n})\}\to 0$ as $n\to\infty$ for any $i\in\{1,\ldots,\ell\}$. Suppose by contradiction that this statement is false. Then there exist $i\in\{1,\ldots,\ell\}$, $\delta>0$, and a subsequence $\{x_{n_{k}}\}$ such that $g_{i}(x_{n_{k}})\geq\delta$.
By the definition of $c_{n}$ (see Step 1 of Algorithm 5) one has $c_{n}\geq n$ for all $n\in\mathbb{N}$. Consequently, if $a<+\infty$, then $c_{n_{k}}g_{i}(x_{n_{k}})\geq a$ and $\varphi(x_{n_{k}},\mu_{n_{k}})=+\infty$ for any sufficiently large $k$, which is impossible.
Let us now consider the case $a=+\infty$. With the use of condition (33) and the fact that the function $s\mapsto P(s,t)$ is nondecreasing one gets
$L_{P}(x_{n_{k}},$ $\lambda_{n_{k}},c_{n_{k}})=f(x_{n_{k}})+\frac{1}{c_{n_{k}}}\sum_{j=1}^{\ell}P(c_{n_{k}}g_{j}(x_{n_{k}}),\lambda_{{n_{k}}j})$
$\geq\inf_{n}f(x_{n})+\frac{1}{c_{n_{k}}}P(c_{n_{k}}g_{i}(x_{n_{k}}),\lambda_{{n_{k}}i})+\frac{1}{c_{n_{k}}}\sum_{j\neq i}P(c_{n_{k}}g_{j}(x_{n_{k}}),\lambda_{{n_{k}}j})$
$\geq\inf_{n}f(x_{n})+\frac{1}{c_{n_{k}}}P(c_{n_{k}}\delta,\lambda_{{n_{k}}i})-\frac{1}{c_{n_{k}}}\sum_{j\neq i}|r(\lambda_{{n_{k}}j})|$
$\geq\inf_{n}f(x_{n})+\frac{1}{c_{n_{k}}}P(c_{n_{k}}\delta,\lambda_{{n_{k}}i})-\frac{1}{n_{k}}.$
By our assumption the sequence $\{f(x_{n})\}$ is bounded below. Hence taking into account the second assumption on the function $P$ one can conclude that $L_{P}(x_{n_{k}},\lambda_{n_{k}},c_{n_{k}})\to+\infty$ as $k\to\infty$, which contradicts the fact that by the weak duality $L_{P}(x_{n},\lambda_{n},c_{n})\leq f_{*}+\varepsilon_{n}$ for any $n\in\mathbb{N}$. Thus, $\max\{0,g_{i}(x_{n})\}\to 0$ as $n\to\infty$ for any $i\in\{1,\ldots,\ell\}$.
Let us finally verify that the sequence $\{(\lambda_{n},c_{n})\}$ satisfies the $\varphi$-convergence to zero condition. Indeed, since $P(0,t)=0$ and the function $s\mapsto P(s,t)$ is continuous and nondecreasing for all $t\in\mathbb{R}_{+}$, for any $n\geq 1$ one can find $t_{n}>0$ such that $c_{n}s<a$ and $P(c_{n}s,\lambda_{ni})<c_{n}/n\ell$ for all $s\leq t_{n}$. Therefore, for any sequence $\{z_{n}\}\subset Q$ satisfying the inequality $\sum_{i=1}^{\ell}\max\{0,g_{i}(z_{n})\}\leq t_{n}$ one has
$\varphi(z_{n},\mu_{n})=\frac{1}{c_{n}}\sum_{i=1}^{\ell}P(c_{n}g_{i}(z_{n}),\lambda_{ni})\leq\frac{1}{n}$
for any $n\in\mathbb{N}$. Consequently, $\limsup_{n\to\infty}\varphi(z_{n},\mu_{n})\leq 0$ and the sequence $\{(\lambda_{n},c_{n})\}$ satisfies the $\varphi$-convergence to zero condition. Therefore, applying Theorem 1 we arrive at the required result. ∎
Remark 8.
The previous theorem significantly generalizes [37, Theorem 5], since it shows that [37, Assumption $(H_{4})$] and the assumptions that
$P(s,0)\geq 0\quad\forall s\in(-\infty,a),\quad\lim_{t\to+\infty}P(s,t)=+\infty\quad\forall s>0$
are completely redundant, while the assumption on the boundedness below of the function $f$ on the set $Q$ (see [37, Assumption 1]) can be replaced by a much less restrictive assumption on the boundedness below of the sequence $\{f(x_{n})\}$.
5 Concluding remarks
We presented the universal convergence theorem for inexact primal-dual methods that (i) can be applied to a wide variety of penalty and augmented Lagrangian methods and (ii) reduces convergence analysis of these methods to verification of some simple conditions related to the convergence of the infeasibility measure to zero. By applying this theorem to convergence analysis of several particular primal-dual methods, we demonstrated that with the use of the universal convergence theorem one can not only easily recover existing results on convergence of various primal-dual methods, but also significantly strengthen and generalize them.
Although the universal convergence theorem is a very general result that can be applied to, perhaps, a majority of existing primal-dual penalty and augmented Lagrangian methods, there are some methods for which one cannot recover existing convergence theorems with the use of the universal convergence theorem. The modified subgradient algorithm from [19] is one such method. One can prove its dual convergence, despite the fact that the infeasibility measure might not converge to zero along the sequences generated by this method (see [5, Example 1]). Moreover, for some closely related primal-dual methods [8, 7, 10] one can prove convergence theorems whose claims are significantly stronger than the claim of the universal convergence theorem (e.g. one can prove strong convergence of the dual variables).
Thus, although the universal convergence theorem unifies and significantly generalizes many existing results on convergence of primal-dual methods, it should be viewed not as a universal result that can completely eliminate the need for detailed convergence analysis, but rather as a useful tool that can significantly simplify this analysis in many important cases.